# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v3

@@ 1.3 Setup
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively  [^4] Note that, the choice of energy functional $E$ is more flexible and corresponds to the ‘gauge fixing’ and is usually dictated by the experimental/physical constraint. Therefore, it does not necessarily have to be an equilibrium energy functional, for instance, a non-equilibrium analogue of energy is used for a non-equilibrium steady-state $E=ss$, a Lyapunov functional or rate functional that generates the steady-state distribution. This threefold orthogonal decomposition of mean dissipation has been highlighted in Ref.[8], and was generalized to a fourfold orthogonal decomposition of mean dissipation in Ref.[5] to systems that break the action-reaction symmetry.. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 8]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 6].
@@ 2.1 Geodesic structure
Equation˜5 quantifies the driving kinetic energy, where the mass is given by the instantaneous ‘local curvature’ of the Lagrangian ($\partial_{{\alpha}}^{2}\mathcal{L}^{*}$). We adopt the shorthand notation $\partial_{{\alpha}}^{2}\mathcal{L}^{*}=\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}$, which satisfies
@@ 2.2 Exact analytical expressions for geodesics
The geodesic is obtained by solving the ODE $\dot{A}_{\alpha}\sqrt{\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}}=v_{qs}$. In addition, we have imposed the constraint $\mathcal{G}(0)=0$, which corresponds to fixing the integration constant. This allows us to compare different geodesics from the same reference point. The exact closed-form analytical expressions for the linear geodesics $\mathcal{G}^{lin}$, close-to-equilibrium geodesics $\mathcal{G}^{cEQ}(A_{\alpha})$ and fEQ geodesics $\mathcal{G}^{fEQ}(A_{\alpha})$ are:
@@ 2.3 Far-from-equilibrium generalizations of
the distance in the geodesic space of the control parameters. This is followed by the definition of $\mathcal{W}_{2}$ and recalling the definition of the geodesic, which physically corresponds to the minimization of distance, and is mathematically equal to the $L^{2}-{norm}$. The definition eq.˜13 of $\mathcal{W}_{2}$ itself is novel, and accounts for the geometric curvature for any control parameter. The existing cEQ optimal transport theory methodology focuses on the control of the state-space. Thus, eq.˜9 leads to equality between the driving EP in ST and the Wasserstein distance in OT,
@@ 2.4 Geometric speed limit for slow driving:
a time-dissipation-geometry tradeoff relation
@@
The quasi-static driving time reveals the fundamental tradeoff between the driving time, dissipation (EP) and geometry for the optimally driven process, and defines the geometric speed limit for the slow-driving process. Therefore, eq.˜15 is a fundamental time-dissipation-geometry tradeoff relation. Importantly, the slow-driving assumption is violated for any $\tau$ and does not necessarily require $\tau\leq\tau_{qs}^{*}$. For the fixed value $\Delta A_{\alpha}=A_{\alpha}^{f}-A_{\alpha}^{i}$ and a fixed available budget for driving EP $\Sigma_{qs}^{*}$, eq.˜11 implies,
@@ 4.1 Multiple control parameters
where eq.˜28 is represented in Einstein’s summation convention, which signifies the summation over the contracted index. $\vec{\lambda}$ is the vector of control parameters. The Hessian of $\mathcal{L}^{*}$ (mass) gives the metric tensor $g_{ij}$,
@@
$\begin{split}\ddot{\lambda}_{i}+\Gamma_{jk}^{i}\dot{\lambda}_{j}\dot{\lambda}_{k}=0,\end{split}$ (30)
@@
By construction, the geodesic $\mathcal{G}(\{\lambda\})$ is the solution that minimizes $\Sigma_{drv}^{qs}$, $\mathcal{G}(\{\lambda\})=\arg\inf_{\{\lambda\}}\left(\Sigma_{drv}^{qs}\right)$. The corresponding optimal protocol is the linear interpolation, which reads,
$\begin{split}\mathcal{G}(\{\lambda\})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(\{\lambda^{i}\})+\frac{t}{\tau}\mathcal{G}(\{\lambda^{f}\}).\end{split}$ (32)
Using eq.˜32, the optimal slow driving EP reads:
@@
Here, $||*||_{2}$ is the $L^{2}-norm$ that quantifies the distance in the Riemannian manifold induced by multiple control parameters. The finite-time optimal control framework developed in section˜3 is based on the MinAP combined with the existence of the slow-driven geodesic $\mathcal{G}(\{\lambda\})$ and incorporating the possibility of kinks. Hence, a multi-parameter finite-time optimal control protocol is rather trivial, provided that a slow driving geodesic eq.˜32 exists and is computed. The corresponding finite-time multi-parameter optimal protocol reads,
$\begin{split}\mathcal{G}_{\tau}(\{\lambda\})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}(\{\lambda^{i}\})+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}(\{\lambda^{f}\}),\end{split}$ (34)
@@ 4.2 Numerical computation of geodesic
We have focused on the exact analytical formulation of the GFTOC framework. However, analytically computing the exact geodesic $\mathcal{G}(\{\lambda\})$ is not always feasible for more sophisticated problems. To this end, a numerical framework is required to compute the geodesic/optimal protocols. Here, we highlight two different numerical algorithms that are implemented to calculate the geodesic. The first algorithm is based on the MinAP, where the geodesic $\mathcal{G}(\{\lambda\})$ is obtained by the minimization of the Lagrangian eq.˜28, and is extended to return the finite-time optimal protocol $\mathcal{G}_{\tau}(\{\lambda\})$, it reads,
@@
Input: $A_{i},A_{f},\tau,g_{ij}$, eqs.˜28, 34 and 35
1:Compute the geodesic $\mathcal{G}(A_{\alpha})$ obtained by the recursive minimization of eq.˜28, and that connects $\mathcal{G}(A_{i})$ and $\mathcal{G}(A_{f})$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
3:Return the optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$.
Note that, Algorithm 1 is modified to directly compute $\mathcal{G}_{\tau}(\{\lambda\})$ instead of $\mathcal{G}(\{\lambda\})$, if the thermodynamic cost for kinks is assigned under the action minimization problem and the endpoint discontinuity in the protocol space is allowed by relaxing the boundary condition constraint. Algorithm 1 extends to the class of numerical optimization methods developed in Ref.[39, 40, 59, 60, 61, 31, 62, 63, 64, 65, 66] and related to machine learning techniques Ref.[67, 68, 69]. However, none of the previous numerical algorithms incorporates kinks as a result of their failure to assign a thermodynamic cost to kinks. To our knowledge, Algorithm 1 is the novel formulation that numerically computes finite-time optimal protocols with kinks and computes the associated finite-time driving EP cost.
A second alternative algorithm computes the finite-time optimal protocols by numerically solving the geodesic eq.˜30 as a boundary value problem. The algorithm for it reads:
@@
Input: $A_{i},A_{f},\tau,g_{ij},\Gamma_{ij}^{k}$, eq.˜30 and eq.˜34
1:Compute the geodesic $\mathcal{G}(A_{\alpha})$ by numerically solving eq.˜30 under the boundary value constraint $A_{i}$ and $A_{f}$ at $t=0$ and $t=\tau$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
3:Return the optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$.
In contrast to Algorithm 1, Algorithm 2 can only compute $\mathcal{G}(A_{\alpha})$, and requires eq.˜34 to compute $\mathcal{G}_{\tau}(A_{\alpha})$. The numerical approach elaborated in this section broadens the applicability of the GFTOC when an exact analytical solution of the geodesic is not available.
@@ 5.1 Optimal control of free energy
The EP defined for the driving of the free energy $\psi_{E}$ by changing the control parameters $\{\lambda\}$ of $E$ reads, $\mathcal{L}_{drv}^{*}=\frac{1}{2}\left(\partial^{2}\psi_{E}/\partial{\lambda_{i}}\partial{\lambda_{j}}\right)\dot{\lambda}_{i}\dot{\lambda}_{j}$. The initial and final control parameter vectors are denoted by $\vec{\lambda}_{E}^{inl}$ and $\vec{\lambda}_{E}^{fnl}$, respectively. The corresponding optimal slow driving EP and the optimal driving protocol read:
$\begin{split}&\Sigma_{qs}^{*}=\frac{1}{2\tau}\left(\mathcal{G}(\vec{\lambda}_{E}^{inl})-\mathcal{G}(\vec{\lambda}_{E}^{fnl})\right)^{2},\\ &\mathcal{G}(\vec{\lambda})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(\vec{\lambda}_{E}^{inl})+\frac{t}{\tau}\mathcal{G}(\vec{\lambda}_{E}^{fnl}).\end{split}$ (36)

### v4

@@ 1.3 Setup
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively  [^4] Note that, the choice of energy functional $E$ is more flexible and corresponds to the ‘gauge fixing’ and is usually dictated by the experimental/physical constraint. Therefore, it does not necessarily have to be an equilibrium energy functional, for instance, a non-equilibrium analogue of energy is used for a non-equilibrium steady-state $E=ss$, a Lyapunov functional or rate functional that generates the steady-state distribution. This threefold orthogonal decomposition of mean dissipation has been highlighted in Ref.[8], and was generalized to a fourfold orthogonal decomposition of mean dissipation in Ref.[5] to non-reciprocal systems that break the action-reaction symmetry.. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 8]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 6]. Throughout this work, we use vector ($\vec{*}$) and equivalent set $\{*\}$ representations interchangeably, depending on the context.
@@ 2.1 Geodesic structure
Equation˜5 quantifies the driving kinetic energy, where the mass is given by the instantaneous ‘local curvature’ of the Lagrangian ($\partial_{{\alpha}}^{2}\mathcal{L}^{*}$). We adopt the shorthand notation $\partial_{{\alpha}}^{2}\mathcal{L}^{*}=\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}$, which satisfies,
@@ 2.2 Exact analytical expressions for geodesics
The geodesic is obtained by solving the ODE $\dot{A}_{\alpha}\sqrt{\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}}=v_{qs}$. In addition, we have imposed the constraint $\mathcal{G}(0)=0$, which corresponds to fixing the integration constant. This allows us to compare different geodesics from the same reference point. The exact closed-form analytical expressions for the linear geodesics $\mathcal{G}^{lin}$, close-to-equilibrium geodesics $\mathcal{G}^{cEQ}(A_{\alpha})$ and far-from-equilibrium geodesics $\mathcal{G}^{fEQ}(A_{\alpha})$ are:
@@ 2.3 Far-from-equilibrium generalizations of
the distance in the geodesic space of the control parameters. This is followed by the definition of $\mathcal{W}_{2}$ and recalling the definition of the geodesic, which physically corresponds to the minimization of distance, and is mathematically equal to the $L^{2}-{norm}$. The definition eq.˜13 of $\mathcal{W}_{2}$ itself is novel, and accounts for the geometric curvature for any control parameter. The existing cEQ optimal transport theory methodology focuses on the control of the state-space. Thus, eq.˜9 leads to equality between the driving EP in ST and the Wasserstein distance in optimal transport theory,
@@ 2.4 Geometric speed limit for slow driving:
time-dissipation-geometry tradeoff relation
@@
The quasi-static driving time reveals the fundamental tradeoff between the driving time, dissipation (EP) and geometry for the optimally driven process, and defines the geometric speed limit for the slow-driving process. Therefore, eq.˜15 is a fundamental ‘time-dissipation-geometry tradeoff relation’. Importantly, the slow-driving assumption is violated for any $\tau$ and does not necessarily require $\tau\leq\tau_{qs}^{*}$. For the fixed value $\Delta A_{\alpha}=A_{\alpha}^{f}-A_{\alpha}^{i}$ and a fixed available budget for driving EP $\Sigma_{qs}^{*}$, eq.˜11 implies,
@@ 4.1 Multiple control parameters
where eq.˜28 is represented in Einstein’s summation convention, which signifies the summation over the contracted index. $\vec{\lambda}$ is the vector of control parameters with $i^{th}$ component $\lambda^{i}$. The Hessian of $\mathcal{L}^{*}$ (mass) gives the metric tensor $g_{ij}$,
@@
$\begin{split}\ddot{\lambda}^{i}+\Gamma_{jk}^{i}\dot{\lambda}^{j}\dot{\lambda}^{k}=0,\end{split}$ (30)
@@
By construction, the geodesic $\mathcal{G}(\vec{\lambda})$ is the solution that minimizes $\Sigma_{drv}^{qs}$, $\mathcal{G}(\vec{\lambda})=\arg\inf_{\{\lambda\}}\left(\Sigma_{drv}^{qs}\right)$. The corresponding optimal protocol is the linear interpolation, which reads,
$\begin{split}\mathcal{G}(\vec{\lambda})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(\vec{\lambda}^{inl})+\frac{t}{\tau}\mathcal{G}(\vec{\lambda}^{fnl}).\end{split}$ (32)
To avoid confusion with the vector index, we denote the vectors of initial and final control parameters by $\vec{\lambda}^{inl}$ and $\vec{\lambda}^{fnl}$, respectively. Using eq.˜32, the optimal slow driving EP reads:
@@
Here, $||*||_{2}$ is the $L^{2}-norm$ that quantifies the distance in the Riemannian manifold induced by multiple control parameters. The finite-time optimal control framework developed in section˜3 is based on the MinAP combined with the existence of the slow-driven geodesic $\mathcal{G}(\vec{\lambda})$ and incorporating the possibility of kinks. Hence, a multi-parameter finite-time optimal control protocol is rather trivial, provided that a slow driving geodesic eq.˜32 exists and is computed. The corresponding finite-time multi-parameter optimal protocol reads,
$\begin{split}\mathcal{G}_{\tau}(\vec{\lambda})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}(\vec{\lambda}^{inl})+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}(\vec{\lambda}^{fnl}),\end{split}$ (34)
@@ 4.2 Numerical computation of geodesic
We have focused on the exact analytical formulation of the GFTOC framework. However, analytically computing the exact geodesic $\mathcal{G}(\vec{\lambda})$ is not always feasible for more sophisticated problems. To this end, a numerical framework is required to compute the geodesic/optimal protocols. Here, we highlight two different numerical algorithms that are implemented to calculate the geodesic.
The first algorithm is based on the MinAP, which focuses on the thermodynamic cost of driving, where the geodesic $\mathcal{G}(\vec{\lambda})$ is obtained by the minimization of the Lagrangian eq.˜28, and is extended to return the finite-time optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$, it reads,
@@
Input: $\vec{\lambda}^{inl},\vec{\lambda}^{fnl},\tau,g_{ij}$, eqs.˜28, 34 and 35
1:Compute the geodesic $\mathcal{G}(\vec{\lambda})$ obtained by the recursive minimization of eq.˜28, and that connects $\mathcal{G}(\vec{\lambda}^{inl})$ and $\mathcal{G}(\vec{\lambda}^{fnl})$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
3:Return the optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$.
Note that, Algorithm 1 is modified to directly compute $\mathcal{G}_{\tau}(\vec{\lambda})$ instead of $\mathcal{G}(\vec{\lambda})$, if the thermodynamic cost for kinks is assigned under the action minimization problem and the endpoint discontinuity in the protocol space is allowed by relaxing the boundary condition constraint. Algorithm 1 extends to the class of numerical optimization methods developed in Ref.[39, 40, 59, 60, 61, 31, 62, 63, 64, 65, 66] and related to machine learning techniques Ref.[67, 68, 69]. However, none of the previous numerical algorithms incorporates kinks as a result of their failure to assign a thermodynamic cost to kinks. To our knowledge, Algorithm 1 is the novel formulation that numerically computes finite-time optimal protocols with kinks and computes the associated finite-time driving EP cost.
A second alternative algorithm computes the finite-time optimal protocols by numerically solving the geodesic eq.˜30 as a boundary value problem, focusing on driving dynamics. The algorithm for it reads:
@@
Input: $\vec{\lambda}^{inl},\vec{\lambda}^{fnl},\tau,g_{ij},\Gamma_{ij}^{k}$, eq.˜30 and eq.˜34
1:Compute the geodesic $\mathcal{G}(\vec{\lambda})$ by numerically solving eq.˜30 under the boundary value constraint $\vec{\lambda}^{inl}$ and $\vec{\lambda}^{fnl}$ at $t=0$ and $t=\tau$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
3:Return the optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$.
In contrast to Algorithm 1, Algorithm 2 can only compute $\mathcal{G}(\vec{\lambda})$, and requires eq.˜34 to compute $\mathcal{G}_{\tau}(\vec{\lambda})$. The numerical approach elaborated in this section broadens the applicability of the GFTOC when an exact analytical solution of the geodesic is not available.
@@ 5.1 Optimal control of free energy
The EP defined for the driving of the free energy $\psi_{E}$ by changing the control parameters $\{\lambda_{E}\}$ of $E$ reads, $\mathcal{L}_{drv}^{*}=\frac{1}{2}\left(\partial^{2}\psi_{E}/\partial{\lambda_{E}^{i}}\partial{\lambda_{E}^{j}}\right)\dot{\lambda}_{E}^{i}\dot{\lambda}_{E}^{j}$. The initial and final control parameter vectors are denoted by $\vec{\lambda}_{E}^{inl}$ and $\vec{\lambda}_{E}^{fnl}$, respectively. The corresponding optimal slow driving EP and the optimal driving protocol read:
$\begin{split}&\Sigma_{qs}^{*}=\frac{1}{2\tau}\left(\mathcal{G}(\vec{\lambda}_{E}^{inl})-\mathcal{G}(\vec{\lambda}_{E}^{fnl})\right)^{2},\\ &\mathcal{G}(\vec{\lambda}_{E})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(\vec{\lambda}_{E}^{inl})+\frac{t}{\tau}\mathcal{G}(\vec{\lambda}_{E}^{fnl}).\end{split}$ (36)